OpenStudy (anonymous):

Medal will be awarded! Find the perimeter of a quadrilateral with vertices at C (−1, 2), D (−2, −1), E (2, −2), and F (1, 1). Round your answer to the nearest hundredth when necessary. 10.92 units 12.68 units 13.43 units 18.27 units

OpenStudy (bradely):

Calculate distances CD,DE,EF, and FC Then add all to get perimeter Formula:sqrt((x2-x1)^2+(y2-y1)^2) Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

okay hold on let me try

OpenStudy (anonymous):

im confused on how to do this actually

OpenStudy (bradely):

C (−1, 2), D (−2, −1) x1=-1,y1=2,x2=-2,y2=-1 CD= sqrt((-2-(-1))^2+(-1-2)^2) =sqrt(1+9) =sqrt10 Same as do remaining and add

OpenStudy (anonymous):

alright give me a second to solve that

OpenStudy (anonymous):

i actually cant do this haha, I'm too confused on how to add them and do all of that sqrt again I'm not taking this class so i wouldn't know.

OpenStudy (imstuck):

I used the distance formula to find the length between each set of points. The distance between C and D, for example, came out to be sqrt10. DF came out to be sqrt17. You use the points in the formula like this: